# DeFine: Enhancing LLM Decision-Making with Factor Profiles and Analogical Reasoning

### v1

@@ Abstract
LLMs are ideal for decision-making due to their ability to reason over long contexts and identify critical factors. However, challenges arise when processing transcripts of spoken speech describing complex scenarios. These transcripts often contain ungrammatical or incomplete sentences, repetitions, hedging, and vagueness. For example, during a company’s earnings call, an executive might project a positive revenue outlook to reassure investors, despite significant uncertainty regarding future earnings. It is crucial for LLMs to incorporate this uncertainty systematically when making decisions. In this paper, we introduce DeFine, a new framework that constructs probabilistic factor profiles from complex scenarios. DeFine then integrates these profiles with analogical reasoning, leveraging insights from similar past experiences to guide LLMs in making critical decisions in novel situations. Our framework separates the tasks of quantifying uncertainty in complex scenarios and incorporating it into LLM decision-making. This approach is particularly useful in fields such as medical consultations, negotiations, and political debates, where making decisions under uncertainty is vital.
@@ 1 Introduction
Large language models are increasingly utilized for decision-making, thanks to their advanced reasoning abilities (Eigner & Händler, 2024). While research has examined various types of reasoning, e.g., deductive, inductive, mathematical, and multi-hop reasoning, most studies have tackled simpler tasks, such as natural language inference and math word problems (Bostrom et al., 2022; Huang & Chang, 2023; Sprague et al., 2024; Mondorf & Plank, 2024). There is a significant gap in handling complex, real-world scenarios, such as making financial investment decisions (Keith & Stent, 2019), where the stakes are high and poor decisions can result in severe consequences. Therefore, it is crucial to understand how LLMs make decisions, allowing domain experts to collaborate with them to make informed, rational decisions in complex situations.
The challenges are compounded when LLMs are required to handle long contexts, extract multiple relevant pieces of information, and make decisions based on this data (Krishna et al., 2023; Laban et al., 2024). Key issues include a tendency to prioritize information at the beginning and end of the context (recency bias; Liu et al. 2023), handling inconsistencies, and mitigating hallucinations in numerical data (Hu et al., 2024a; b). Current tools such as chain-of-thought (CoT; Wei et al. 2023), tree-of-thought (ToT; Yao et al. 2023), Reflexion (Shinn et al., 2023), are designed to provide reasoning traces for LLM decisions; however, their explanations remain ambiguous. They lack precise, quantitative insights into key factors and the degree of uncertainty involved. Consequently, decision-makers are left with doubts about the reliability of these decisions and how they can be improved. There is a pressing need to enhance the verifiability of LLMs in complex decision-making scenarios to ensure their dependability and effectiveness.
We present DeFine, a new framework designed to build probabilistic factor profiles from transcripts of spoken speech that describe complex scenarios. These transcripts are often excessively long, containing ungrammatical sentences, repetitions, hedging, and vagueness (Sawhney et al., 2020; Medya et al., 2022). For example, during a quarterly earnings call, a company executive might project a positive revenue outlook to boost investor confidence, despite significant uncertainties surrounding these projections (Mukherjee et al., 2022). DeFine constructs a factor profile for each transcript that summarizes essential information into a set of factors and estimates the probabilities of potential outcomes for these factors. Moreover, we employ the Bradley-Terry model (Bradley & Terry, 1952) to identify dominant factors and evaluate how these factors collectively impact decision-making. Our factor profiles are designed to capture the nuances in spoken transcripts. This includes not just what is explicitly stated, but also the implications of details that are omitted, providing a viable method for quantifying uncertainty.
Our research integrates probabilistic factor profiles with analogical reasoning, a type of reasoning that identifies connections between similar situations to facilitate knowledge transfer from a familiar context to a novel one (Webb et al., 2023; Yasunaga et al., 2024). It helps decision-makers draw parallels between current situations and past experiences, thus effectively leveraging historical insights to inform their decisions. Instead of relying on traditional text matching, we use factor profiles to retrieve analogous examples, which identifies historical cases with similar levels of uncertainty across key dimensions. Analogical reasoning further sets our work apart from traditional Bayesian inference frameworks used in decision-making (Halawi et al., 2024; Lin et al., 2024; Liu et al., 2024). The latter often require extensive sampling during inference, which tends to increase inference costs and potentially leads to latency issues. The contributions of our research are summarized as follows.
  • We introduce DeFine, a new framework designed to enhance decision-making in LLMs. DeFine utilizes probabilistic factor profiles to quantify uncertainty in complex scenarios, along with analogical reasoning that leverages insights from similar past experiences to guide LLMs in making crucial decisions in novel situations. Our framework aims to boost the utility of LLMs by adding a layer of transparency to their decision-making processes.
  • Our research has produced actionable insights for predicting stock movement trends by analyzing earnings call transcripts. These insights enable investors to make informed, data-driven decisions. Furthermore, our approach has potential applications beyond finance, including in fields such as medical consultations and political debates (Lehman et al., 2022), where the discussions involve complex issues and the decisions made can have significant consequences. [^1] We plan to make the data and code publicly available upon acceptance to facilitate research in this area.
@@
2 Overview and Factor Analysis
@@
LLMs have shown great promise in financial services (Reddy et al., 2024; Lee et al., 2024), yet the specifics of their decision-making processes remain largely unexplored. In this paper, we investigate how LLMs can guide investment decisions by analyzing earnings call transcripts. The transcripts contain a rich mix of textual and numerical data, presenting a unique challenge for LLMs in making rational decisions. An earnings call is a teleconference in which the management of a public company discusses its financial results with analysts and investors for a specific reporting period, such as a quarter or a fiscal year. [^2] https://en.wikipedia.org/wiki/Earnings_call The transcript generally consists of two sections: the initial prepared remarks from the company’s executives, and a subsequent Q&A session. During these calls, executives provide a deep dive into the company’s financials, discuss key performance indicators, and share strategic plans for the future. An excerpt from a typical earnings call transcript is shown in Figure 1.
@@
Our goal is to provide investment recommendations based on earnings call transcripts (ECTs), using a five-tier system: Strong Buy, Buy, Hold, Sell, and Strong Sell. We choose this approach over a simple binary classification (Ni et al., 2024) to give a clearer and more nuanced assessment of the investment opportunity. We highlight the key drivers in decision-making by identifying a small set of factors from a lengthy transcript (Eigner & Händler, 2024; Feng et al., 2024). For example, in an earnings call, discussions of financials such as revenue, expenses, and profit margins can be overwhelming. A factor profile helps to distill these discussions into multiple variables, effectively reducing information redundancy and allowing decision-makers to focus on the most impactful factors. Crucially, a factor profile offers a comprehensive view on an earnings call. If critical elements such as debt levels are not addressed by company executives, they can be marked as ‘unknown or uncertain.’ This contrasts with textual summaries of the transcript (Cho et al., 2021; Khatuya et al., 2024), which may be biased toward the topics emphasized by executives and discussed during the Q&A session.
2.1 Factor Profile
[…]
@@ 2.1 Factor Profile
In this study, we focus on a curated set of 15 factors, categorized into three groups: macroeconomic influences (e.g., economic health, market sentiment), company-specific dynamics (e.g., mergers and major acquisitions, product launches), and historical financial metrics (e.g., past earnings, stock prices). These factors were carefully selected through an iterative process of querying the LLM for key variables crucial in forecasting stock movements following earnings announcements. We intentionally limited our variable set to 15 factors, each with two to three potential outcomes, as detailed below. By distilling the analysis to a few significant predictors, our approach balances complexity and performance, while also allowing for future integration of domain-specific factors identified by financial analysts.
@@ 2.2 Analyzing Key Factors Using the Bradley-Terry Model
The Bradley-Terry model is a probabilistic framework used for estimating the relative strengths of items based on pairwise comparisons, and the outcome of each comparison indicates which of the two items is ‘better’ in a specific context (Bradley & Terry, 1952). This model has been widely used for ranking purposes in sports tournaments, LLM preference studies, and other domains where pairwise comparison data is available (Hu et al., 2023; Zhu et al., 2024). In this model, we estimate parameters that represent the strength of each factor. These parameters are generally presented on a logistic scale, where the probability that factor A is considered more significant than factor B is modeled as:
@@ 4 Analogical Reasoning
5 Data Collection
Our dataset contains 11,950 earnings call transcripts from S&P 500 and NASDAQ 500 companies, gathered from the Motley Fool over the period of 2017–2024. The Motley Fool is a well-regarded financial service website that regularly publishes earnings call transcripts from U.S. companies. We make sure to follow their terms of use carefully during data collection. We do not use audio recordings or analyze acoustic or prosodic features. Each transcript is formatted as a JSON object, including the company’s stock ticker, the date of the earnings announcement, participant names and their affiliations, executive prepared remarks, and a series of question-answer pairs from the Q&A session. Table 2 presents the statistics of our dataset. Each transcript averages 10,187 tokens and 133 sentences. They are sourced from 869 companies, each contributing an average of 14 transcripts. We obtain company stock prices from Yahoo Finance via the yfinance package and financial metrics such as revenue and earnings per share (EPS) from Alpha Advantage. Our dataset spans from 2017 to 2024. It enhances previous studies which examined earnings call transcripts from 2002–2010 (Li et al., 2020); these earlier transcripts may already be used in LLM pretraining. To avoid data contamination, we established a new test set consisting of the most recent 587 transcripts from 2024, which are beyond the pretraining cut-off date for LLMs.
6 Experiments
@@ 6.1 Decision-Making with DeFine
We test our system, DeFine, against different decision-making strategies: (a) LLM+CoT+Trans: We feed the entire earnings call transcript to the LLM and then use the chain-of-thought to assign a label, with both labels and their interpretations also provided to the LLM. (b) LLM+CoT+Summ and LLM+CoT+Factors: These approaches use a summarize-then-predict strategy. LLM+CoT+Summ simplifies the transcript into a textual summary, while LLM+CoT+Factors condenses it into a factor profile. It is considered a structured summary, unlike the textual summary produced by direct LLM prompting. Details on the prompts used for these methods can be found in the Appendix.
@@ 6.2 Influential Factors
Table 4: Among the three variants, DeFine-BT-Cross Sector achieves the highest scores, suggesting that considering pairwise comparisons from a diverse range of companies can enhance the predictions of stock movements.
Table 5: Influential factors and outcomes that drive bullish investment decisions in the Consumer Defensive sector, such as food and beverage, household products, and grocery stores.
Table 6: Factors and outcomes that drive bullish investment decisions in the Technology sector, including industry leaders such as Apple, Microsoft, Amazon, Google, and Meta.
Tables 6 and 6 highlight influential factors impacting investment decisions in the Consumer Defensive and Technology sectors, as identified by the Bradley-Terry model. In Consumer Defensive, which includes industries like food and beverage, household products, and grocery stores, significant drivers are natural disasters and black swan events, political events and government policies, and geopolitical issues. These challenging macroeconomic circumstances often lead to buy-in decisions from investors. In contrast, the Technology sector, with industry leaders such as Apple, Microsoft, Amazon, Google, Meta, and Nvidia, shows that decisions to invest often hinge on unclear or uncertain factors. Technology stocks have seen considerable growth from 2017–2024. This pattern suggests that investment models may favor purchases in these companies despite encountering negative issues in earnings announcements.
@@ 6.3 Insights into Analogical Reasoning
Figure 3: Analogical reasoning utilizes a number of analogous examples to inform decision-making in LLMs. We adjust $K$ from 3 to 9 and observe its impact on the F-Score.
@@ 7 Related Work
Analogical Reasoning. This type of reasoning identifies connections between similar, though not identical, situations to transfer knowledge from a known context to a new one (Webb et al., 2023; Ozturkler et al., 2023; Yu et al., 2024; Yuan et al., 2024; Sourati et al., 2024; Yasunaga et al., 2024). It helps decision makers draw parallels between current situations and past experiences, effectively leveraging historical insights. Analogical reasoning plays a crucial role in various fields, e.g., doctors apply knowledge from one disease to diagnose another, and lawyers use past rulings to argue new cases (Lehman et al., 2022; Charmet et al., 2022; Cao et al., 2024a). This ability to recognize and use similarities in different situations is important for decision-making.
@@
While zero-shot analogical reasoning is a desired capability for LLMs, recent studies show they lack the robustness and generality of human analogy-making, as evidenced by counterexamples in tasks such as letter string analogies (Hodel & West, 2024; Lewis & Mitchell, 2024). Musker et al. (2024) test both humans and LLMs on tasks that require transferring semantic structure and content between domains. Yasunaga et al. (2024) introduce analogical prompting, where LLMs self-generate relevant examples using prompts such as “# Recall relevant problems and solutions:” before solving the original problem; Qin et al. (2024) find that the accuracy of self-generated examples is key to eliciting such capability. Unlike previous research, our study employs probabilistic factor profiles to model analogical reasoning, grounding our approach in solid mathematical principles.
@@
LLM Decision-Making under Uncertainty. The use of LLMs in decision-making has surged due to their remarkable ability to reason over complex scenarios (Halawi et al., 2024; Lin et al., 2024; Ye et al., 2024; Band et al., 2024). However, the challenge of balancing a multitude of often conflicting factors in decision making remains understudied. For example, Falck et al. (2024) investigate whether adding more data points in in-context learning reduces uncertainty, as typically expected in Bayesian learning, and find evidence against this theory. The DeLLMa framework (Liu et al., 2024) incorporates uncertainty into LLM decision-making using Bayesian networks and has been tested on tasks such as agriculture planning and finance. Feng et al. (2024) employ LLM entailment to map factors to context and utilize trained Bayesian models for probability estimation. Our work builds on these initiatives by integrating analogical reasoning with factor profiles to enhance the accuracy and transparency of LLM decision-making.
@@
Financial Forecasting. Recent advancements in LLMs have revolutionized traditional financial tasks (Keith & Stent, 2019; Sawhney et al., 2020; 2021; Chuang & Yang, 2022; Ang & Lim, 2022; Sang & Bao, 2022; Medya et al., 2022; Wang et al., 2023; Koa et al., 2024; Srivastava et al., 2024). Notably, Chen et al. (2022) introduce FinQA, a dataset constructed from financial statements for assessing LLMs’ multi-step numerical reasoning. Moreover, TAT-QA (Zhu et al., 2021) tackles QA over tabular and textual data; FiNER (Loukas et al., 2022) focuses on numerical entity recognition; DocFinQA (Reddy et al., 2024) is a dataset designed for long-document financial QA; RiskLabs (Cao et al., 2024b) employs LLMs for financial risk assessments. Nie et al. (2024) provide a comprehensive survey on the use of LLMs across various financial domains. Our study focuses on analyzing earnings transcripts to understand how LLMs handle the ambiguities inherent in spoken language, thus providing insight into their decision-making under uncertainty. The research findings have broader applications including medical consultations, negotiations, and political debates.

### v2

# DeFine: Decision-Making with Analogical Reasoning over Factor Profiles

@@ Abstract
LLMs are ideal for decision-making thanks to their ability to reason over long contexts. However, challenges arise when processing speech transcripts that describe complex scenarios, as they are verbose and include repetition, hedging, and vagueness. E.g., during a company’s earnings call, an executive might project a positive revenue outlook to reassure investors, despite uncertainty regarding future earnings. It is crucial for LLMs to incorporate this uncertainty systematically when making decisions. In this paper, we introduce DeFine, a modular framework that constructs probabilistic factor profiles from complex scenarios. It then integrates these profiles with analogical reasoning, leveraging insights from similar past experiences to guide LLMs in making critical decisions in new situations. Our framework separates the tasks of quantifying uncertainty and incorporating it into LLM decision-making. This approach is particularly useful in areas such as consulting and financial deliberation, where making decisions under uncertainty is vital.
[…]
tcb@breakable
DeFine: Decision-Making with Analogical Reasoning over Factor Profiles
Yebowen Hu, Xiaoyang Wang, Wenlin Yao, Yiming Lu,§ Daoan Zhang Hassan Foroosh, Dong Yu, Fei Liu§ University of Central Florida   Tencent AI Lab, Seattle University of Rochester   §Emory University {yebowen.hu, hassan.foroosh}@ucf.edu    daoan.zhang@rochester.edu {shawnxywang, wenlinyao, dyu}@global.tencent.com  fei.liu@emory.edu
@@ 1 Introduction
Large language models are increasingly relied on for decision-making, thanks to their powerful reasoning capabilities OpenAI et al. (2024); Anthropic (2025); Kavukcuoglu (2025). While research has advanced rapidly in areas such as math, coding, and logical reasoning Bostrom et al. (2022); Huang and Chang (2023); Sprague et al. (2024); Mondorf and Plank (2024); Li et al. (2025); Ren et al. (2025), there is growing interest in exploring how LLMs reason through complex, real-world environments to make high-stakes decisions, such as financial investments (Peters, 2024).
The challenges are compounded when processing long contexts (Krishna et al., 2023; Laban et al., 2024). Core issues may include recency bias, hallucinations, numerical inconsistencies, and more (Liu et al., 2023; Hu et al., 2024a, b; Gao et al., 2024). While these models are designed to produce reasoning traces, their explanations can remain vague and even unfaithful Chen et al. (2025). Critically, LLMs still lack precise, quantitative insight into the key factors that lead to their final decisions, as well as mechanisms for incorporating uncertainty into their decision-making.
In this paper, we present DeFine, a new framework for constructing probabilistic factor profiles from speech transcripts that describe complex scenarios, and leveraging these profiles to enhance decision-making. For example, during an earnings call, a company executive might project strong revenue growth to boost investor confidence, despite substantial uncertainty surrounding these projections (Mukherjee et al., 2022). As illustrated in Figure 1, DeFine generates a structured factor profile from each transcript, capturing not only what is explicitly stated, but also the implications of what is left unsaid. It then uses the BT model (Bradley and Terry, 1952) to identify dominant factors and evaluate how these factors collectively impact decision-making.
Our research integrates probabilistic factor profiles with analogical reasoning, a type of reasoning that identifies connections between similar situations to facilitate knowledge transfer from a familiar context to a new situation (Webb et al., 2023; Yasunaga et al., 2024). Instead of relying on text matching, we use factor profiles to retrieve analogous examples, which are historical cases with similar levels of uncertainty across key dimensions. Analogical reasoning further sets our work apart from traditional Bayesian inference models used in decision-making (Halawi et al., 2024; Lin et al., 2024; Liu et al., 2024), which usually lack direct connections to historical cases. Our key contributions are as follows.
  • We introduce DeFine, a modular framework that enhances LLM decision-making in complex scenarios. It transforms lengthy speech transcripts into structured factor profiles, which contain key decision factors and their associated uncertainties. Analogical reasoning retrieves comparable profiles and draws on similar past cases to make critical decisions. DeFine thereby adds a layer of transparency into the decision process.
  • Our research demonstrates how to effectively extract decision factors from lengthy transcripts and use them to forecast post-earnings stock movements. This approach has the potential to generalize to other domains, such as consulting, financial investments, and political debates (Lehman et al., 2022), where discussions are complex and decisions carry significant consequences.
2 The DeFine Framework
@@
We investigate how LLMs make critical decisions about post-earnings stock movements using earnings call transcripts (Ni et al., 2024; Peters, 2024). An earnings call is a teleconference in which company executives discuss financial results with analysts and investors for a given quarter or fiscal year. As shown in Figure 5 in the Appendix, these transcripts typically consist of two parts: prepared remarks from the company’s executives and a subsequent Q&A session. During these calls, the executives provide a deep dive into the company’s financials, discuss key performance indicators, and share strategic plans for the future.
@@
However, discussions of financials, e.g., revenue, expenses, and profit margins, can be overwhelming. A factor profile seeks to distills these discussions into multiple variables, allowing decision-makers to focus on the most impactful factors (Eigner and Händler, 2024; Feng et al., 2024). Notably, if critical elements such as debt levels are not addressed by company executives, they can be marked as “unknown or uncertain.” This approach contrasts with traditional textual summaries of the transcript (Cho et al., 2021, 2022; Khatuya et al., 2024), which may be biased toward the topics emphasized by executives and discussed during the Q&A session.
2.1 Constructing Factor Profiles
Let $X$ be an earnings call transcript used to predict a stock investment decision $Y$ with five outcomes: strong buy, buy, hold, sell, and strong sell. We construct a factor profile for each transcript, defining a set of factors =$\mathcal{F}=\{F_{1},F_{2},\dots,F_{n}\}$, where each factor $F_{i}$ has potential outcomes $\{O_{i1},O_{i2},\dots,O_{im}\}$. The likelihood of each outcome, given $X$, is modeled by a probabilistic function $P(O_{ij}|X)$. These probabilities are inferred using a methodology that optimally integrates textual reasoning with quantitative analysis. Thus, each factor outcome’s probability informs the aggregation model that predicts $Y$.
@@ 2.1 Constructing Factor Profiles
Our study focuses on 15 key factors grouped into three categories: macroeconomic influences (e.g., economic health, market sentiment), company-specific dynamics (e.g., mergers and major acquisitions, product launches), and historical financial metrics (e.g., past earnings, stock prices). These factors were iteratively selected by querying the LLM for key variables in forecasting stock movements. We limit the set to 15 factors, each with two to three possible outcomes, ensuring a balance between complexity and performance while allowing future integration of analyst-identified factors. The full list of factors is in the Appendix.
@@ 2.2 Analyzing Key Factors Using the Bradley-Terry Model
The Bradley-Terry model is a probabilistic framework used for estimating the relative strengths of items based on pairwise comparisons, and the outcome of each comparison indicates which of the two items is ‘better’ in a specific context (Bradley and Terry, 1952). This model has been widely used for ranking purposes in sports tournaments, LLM preference studies, and other domains where pairwise comparison data is available (Hu et al., 2023; Zhu et al., 2024). In this model, we estimate parameters that represent the strength of each factor. These parameters are generally presented on a logistic scale, where the probability that factor A is considered more significant than factor B is modeled as:
@@ 4 Analogical Reasoning
5 Data Collection
Our dataset contains 11,950 earnings call transcripts from S&P 500 and NASDAQ 500 companies, gathered from the Motley Fool over the period of 2017–2024. The Motley Fool is a well-regarded financial service website that regularly publishes earnings call transcripts from U.S. companies. We make sure to follow their terms of use carefully during data collection. We do not use audio recordings or analyze acoustic or prosodic features. Each transcript is formatted as a JSON object, including the company’s stock ticker, the date of the earnings announcement, participant names and their affiliations, executive prepared remarks, and a series of question-answer pairs from the Q&A session.
Table 2 presents the statistics of our dataset. Each transcript averages 10,187 tokens and 133 sentences. They are sourced from 869 companies, each contributing an average of 14 transcripts. We obtain company stock prices from Yahoo Finance via the yfinance package and financial metrics such as revenue and earnings per share (EPS) from Alpha Advantage. Our dataset spans from 2017 to 2024. It enhances previous studies which examined earnings call transcripts from 2002–2010 (Li et al., 2020); these earlier transcripts may already be used in LLM pretraining. To avoid data contamination, we established a new test set consisting of the most recent 587 transcripts from 2024, which are beyond the pretraining cut-off date for LLMs.
6 Experiments
Table 4: Influential factors that drive bullish investment decisions in the Consumer Defensive sector, e.g., food and beverage, household products, and grocery stores.
Table 5: Factors that drive bullish investment decisions in the Technology sector, including industry leaders such as Apple, Microsoft, Amazon, Google, and Meta.
@@ 6.1 Decision-Making with DeFine
We evaluate our system, DeFine, against various decision-making strategies: (a) LLM+CoT+Trans: The LLM processes the full earnings call transcript, using chain-of-thought reasoning to assign a label with interpretations. (b) LLM+CoT+Summ and LLM+CoT+Factors: Both follow a summarize-then-predict approach. LLM+CoT+Summ generates a textual summary, while LLM+CoT+Factors condenses information into a structured factor profile. Details on the prompts are in the Appendix.
@@ 6.2 Influential Factors
Table 7: DeFine-BT-Cross Sector achieves the highest scores, suggesting that considering pairwise comparisons from a diverse range of companies can enhance the predictions of stock movements.
Tables 5 and 5 highlight influential factors impacting investment decisions in the Consumer Defensive and Technology sectors, as identified by the Bradley-Terry model. In Consumer Defensive, which includes industries like food and beverage, household products, and grocery stores, significant drivers are natural disasters and black swan events, political events and government policies, and geopolitical issues. These challenging macroeconomic circumstances often lead to buy-in decisions from investors. In contrast, the Technology sector, with industry leaders such as Apple, Microsoft, Amazon, Google, Meta, and Nvidia, shows that decisions to invest often hinge on unclear or uncertain factors. Technology stocks have seen considerable growth from 2017–2024. This pattern suggests that investment models may favor purchases in these companies despite encountering negative issues in earnings announcements.
@@ 6.3 Insights into Analogical Reasoning
Figure 4: Analogical reasoning in LLMs uses analogous examples for decision-making. We vary $K$ from 3 to 9 to assess its impact on the F-Score.
@@ 7 Related Work
. This type of reasoning identifies connections between similar, though not identical, situations to transfer knowledge from a known context to a new one (Webb et al., 2023; Ozturkler et al., 2023; Yu et al., 2024; Yuan et al., 2024; Sourati et al., 2024; Yasunaga et al., 2024). It helps decision makers draw parallels between current situations and past experiences, effectively leveraging historical insights. Analogical reasoning plays a crucial role in various fields, e.g., doctors apply knowledge from one disease to diagnose another, and lawyers use past rulings to argue new cases (Lehman et al., 2022; Charmet et al., 2022; Cao et al., 2024a). This ability to recognize and use similarities in different situations is important for decision-making.
@@
While zero-shot analogical reasoning is a desired capability for LLMs, recent studies show they lack the robustness and generality of human analogy-making, as evidenced by counterexamples in tasks such as letter string analogies (Hodel and West, 2024; Lewis and Mitchell, 2024). Musker et al. (2024) test both humans and LLMs on tasks that require transferring semantic structure and content between domains. Yasunaga et al. (2024) introduce analogical prompting, where LLMs self-generate relevant examples using prompts such as “# Recall relevant problems and solutions:” before solving the original problem; Qin et al. (2024) find that the accuracy of self-generated examples is key to eliciting such capability. Unlike previous research, our study employs probabilistic factor profiles to model analogical reasoning, grounding our approach in solid mathematical principles.
Decision-Making under Uncertainty.
@@
The use of LLMs in decision-making has surged due to their remarkable ability to reason over complex scenarios (Halawi et al., 2024; Lin et al., 2024; Ye et al., 2024; Band et al., 2024). However, the challenge of balancing a multitude of often conflicting factors in decision making remains understudied. For example, Falck et al. (2024) investigate whether adding more data points in in-context learning reduces uncertainty, as typically expected in Bayesian learning, and find evidence against this theory. The DeLLMa framework (Liu et al., 2024) incorporates uncertainty into LLM decision-making using Bayesian networks and has been tested on tasks such as agriculture planning and finance. Feng et al. (2024) employ LLM entailment to map factors to context and utilize trained Bayesian models for probability estimation. Our work builds on these initiatives by integrating analogical reasoning with factor profiles to enhance the accuracy and transparency of LLM decision-making.
Financial Forecasting.
@@
Recent advancements in LLMs have revolutionized traditional financial tasks (Keith and Stent, 2019; Sawhney et al., 2020, 2021; Chuang and Yang, 2022; Ang and Lim, 2022; Sang and Bao, 2022; Medya et al., 2022; Wang et al., 2023; Koa et al., 2024; Srivastava et al., 2024). Notably, Chen et al. (2022) introduce FinQA, a dataset constructed from financial statements for assessing LLMs’ multi-step numerical reasoning. Moreover, TAT-QA (Zhu et al., 2021) tackles QA over tabular and textual data; FiNER (Loukas et al., 2022) focuses on numerical entity recognition; DocFinQA (Reddy et al., 2024) is a dataset designed for long-document financial QA; RiskLabs (Cao et al., 2024b) employs LLMs for financial risk assessments. Nie et al. (2024) provide a comprehensive survey on the use of LLMs across various financial domains. Our study focuses on analyzing earnings transcripts to understand how LLMs handle the ambiguities inherent in spoken language, thus providing insight into their decision-making under uncertainty. The research findings have broader applications including medical consultations, negotiations, and political debates.